A 100% concentration is to be mixed with a mixture having a concentration of 42% to obtain 53 gallons of a mixture with a concentration of 71%. How much of the 100% concentrate will be needed?
Answer by macston(5194)

.
x=amount of 100% solution
.
1.00x+0.42(53-x)=0.71(53)
1.00x+22.26-0.42x=37.63
0.58x=15.37
x=26.5
ANSWER: 26.5 gallons of 100% solution will be needed.
.
CHECK:
53gal-x=53gal-26.5gal=26.5gal
26.5 gallons of 42% solution will be needed.
.
1.00(26.5)+0.42(26.5)=0.71(53)
26.5+11.13=37.63
37.63=37.63